Home‎ > ‎Crew 133‎ > ‎

109-GreenHabReport

XO Gartrelle to Cap Com

The Greenhab is operating normally.

The fan is operating normally set at 68 degrees F.

The heater was set to 40 degrees F at 9AM
The heater was set to 75 degrees F at 5PM

All plants except for the cactus in the Zen Room received appropriate amounts of water today. 
The onion, basil and cilantro seedlings have not germinated. The Mars Shoot is in good shape and has grown just slightly. The hydroponic vegetable seedlings are beginning to sprout, particularly the lettuce. We may see above ground sprouts in the next day or so. The hydroponic seedlings were placed in the green hab at 9AM returned to the upper Main hab for the night at 5PM. 

Tomorrow we will finalize our plans for starting the hydroponic operations on Sunday Jan 12. We will follow the excellent template left by Crew 132.

Photos from taken in the Greenhab today are attached. 

Greenhab readings for today are as follows:

                               Current                 Max                  Min   
Temperature         68 Degrees          82 Degrees       41 Degrees
Humidity                   17%                    21%                   14%